# Strong-to-weak spontaneous breaking of 1-form symmetry and intrinsically mixed topological order

### v2

@@ II.2.2 Relation to the Choi state
The last line can be expressed as a sum of expectation values in the unnormalized Choi state $|\tilde{\rho}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩.
@@
$|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\frac{1}{\sqrt{\mathrm{Tr}(\rho^{2})}}\sum_{n}p_{n}|\psi_{n,u}\rangle\otimes|\psi_{n,l}^{*}\rangle,$ (14)
The unnormalized Choi state $|\tilde{\rho}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ is the same as $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ except with norm equal to the purity of $\rho$. If $\rho$ is a pure state, then $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ is a tensor product: $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|\psi_{u}\rangle|\psi_{l}^{*}\rangle$. Therefore, a pure topologically ordered density matrix maps onto the topological order stacked with its time reversal partner. If $\rho$ is not pure, then there is generally some entanglement between spins in the upper and lower Hilbert spaces.
@@
$\mathrm{Tr}(\mathrm{Tr}_{C}(\rho)^{2})=\frac{1}{D_{C}}\sum_{\sigma_{n}\in C}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\tilde{\rho}|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\tilde{\rho}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$⟩ , (15)
@@
$\begin{split}&I_{2}(A:C|B)\\ &=-\log\frac{\sum_{\sigma_{n}\in C,\sigma_{m}\in A}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}{\sum_{\sigma_{n}\in C,\sigma_{m}\in A}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}\end{split}$ (16)
where we simplified the expression using the normalized Choi state $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=\frac{1}{\sqrt{\mathrm{Tr}(\rho^{2})}}|\tilde{\rho}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩. (LABEL:I2choi) is the main result of this section. It connects the Rényi-2 Markov length to connected correlation functions in the Choi state.
@@
$\begin{split}\Delta_{n,m}&=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\\ &-\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\end{split}$ (17)
@@
$\begin{split}&I_{2}(A:C|B)\\ &=-\log\left(1-\frac{\sum_{m,n}\Delta_{n,m}}{\sum_{\sigma_{n}\in C,\sigma_{m}\in A}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}\right)\\ &\sim\frac{\sum_{m,n}\Delta_{n,m}}{\sum_{\sigma_{n}\in C,\sigma_{m}\in A}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}\end{split}$ (18)
@@
where we expanded to leading order in $\frac{\sum_{m,n}\Delta_{n,m}}{\sum_{\sigma_{n}\in C,\sigma_{m}\in A}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}}$.
@@
Suppose that $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ displays exponential decay of connected correlations. Then $\Delta_{n,m}=0$ if $\sigma_{n,u}\sigma_{n,l}^{\dagger}=\mathbf{1}$ or $\sigma_{m,u}\sigma_{m,l}^{\dagger}=\mathbf{1}$ and $\Delta_{n,m}$ decays exponentially in $r$ otherwise. Therefore, $I_{2}(A:C|B)$ seems to also decays exponentially in $r$, and we obtain a finite Rényi-2 Markov length.
@@
For example, if $\rho=|\psi\rangle\langle\psi|$ is a pure state so $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|\psi\rangle|\psi\rangle$, then assuming that $\rho$ can be written as a product of $|A\cup B\cup C|$ local projectors  [^13] Note that if $|\psi\rangle$ is related by a local unitary $U$ to a fixed point state $\prod_{i}P_{i}$ where each $P_{i}$ is a local projector, then $|\psi\rangle$ can also be written as a product of local projectors: $|\psi\rangle\langle\psi|=\prod_{i}U^{\dagger}P_{i}U$., the denominator of (LABEL:i2expand) is of order $2^{|A|+|C|}$. On the other hand, the numerator is bounded (very loosely) by $(2^{|A|+|C|}-2^{|A|}-2^{|C|})e^{-\mu r}$ for some constant $\mu$, so $I_{2}(A:C|B)$ decays exponentially with $r$. [Edits] More generally when $\rho$ is mixed, then the denominator will typically be of order $e^{\nu(|A|+|C|)}$ for some $\nu$ (for example, if $\rho$ is the completely mixed state, $e^{\nu}=4$) and the numerator is loosely upper bounded by $(e^{\nu(|A|+|C|)}-e^{\nu|A|}-e^{\nu|C|})e^{-\mu r}$ so $I_{2}(A:C|B)$ is upper bounded by a function exponentially decaying in $r$. If instead $\Delta_{n,m}$ decays with a power law for $\mathcal{O}(e^{\nu(|A|+|C|)})$ operators, then $I_{2}(A:C|B)$ cannot be fit to an exponential decay. Attempting to fit $I_{2}(A:C|B)$ to exponential decay in $r$ would lead to a diverging length scale, which is the diverging Rényi-2 Markov length.
@@
Therefore, we expect that the Rényi-2 Markov length diverges whenever the Choi state encounters phase transitions where operators of the form $O_{u}O_{l}^{\dagger}$ have power-law decay of correlations. Note that since the expression (LABEL:I2choi) only has operators that are diagonal between the upper and lower Hilbert spaces, it is not bounded by decay of correlations of operators solely supported in a single Hilbert space, i.e. those of the form $\rho|\sigma_{n,u}\sigma_{m,u}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}⟨$⟩. However, due to the $\mathrm{SWAP}^{*}$ symmetry of the Choi state (explained further in Sec. II.4), the above scenario where there is only a diverging correlation length in $\mathcal{H}_{u}$ or $\mathcal{H}_{l}$ is ruled out.
@@
In the example of decohered toric code Fan et al. (2024); Bao et al. (2023); Lee et al. (2023); Zou et al. (2023); Sang and Hsieh (2024) (see Appendix D for a review) where we put the state $|tc\rangle\langle tc|$ through an incoherent $Z$ noise channel (4) and $X$ noise channel (same as (4) except with $Z_{e}$ replaced by $X_{e}$), the Choi state is initially described by two copies of toric code $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|tc\rangle|tc\rangle$. Under decoherence, the Choi state generically demonstrates two transitions, because there is a transition from two copies of toric code to a single copy, and then from a single copy to the trivial state. In Appendix D.1, we show by explicit calculation in this example that $I_{2}(A:C|B)$ is related to the free energy of a point defect in the 2D classical Ising model, and $\xi_{2}$ diverges at two points.
@@ II.3 Strong and weak 1-form symmetries and order/disorder parameters
$U_{u}U_{l}^{T}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ (21)
where $u,l$ label the upper and lower Hilbert spaces corresponding to the ket and bra, respectively. The order parameter can be either two point correlator in the upper Hilbert space $\rho|Z_{u,i}Z_{u,j}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}∝ ⟨$⟩ or in the lower Hilbert space $\rho|Z_{l,i}Z_{l,j}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}∝ ⟨$⟩, because neither $Z_{u,i}$ nor $Z_{u,j}$ commutes with $U_{u}U_{l}^{T}$, but both $Z_{u,i}Z_{u,j}$ and $Z_{l,i}Z_{l,j}$ commute with $U_{u}U_{l}^{T}$. Since only operators from one of the two Hilbert spaces is involved, we label the order parameter $O_{1}$ with subscript $1$. We cannot use $O_{2}$ as an order parameter because $Z_{u,i}Z_{l,i}$ is not charged under $U_{u}U_{l}^{T}$. By comparison, the disorder parameter is the symmetry action restricted to a subsystem (a $d$-ball in $d$ spatial dimensions). Since the weak symmetry acts on both Hilbert spaces, the disorder parameter lives on both Hilbert spaces as well, which is $\rho|\prod_{i\in A}X_{u,i}X_{l,i}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}∝ ⟨$⟩, which explains the subscript $2$ in $D_{2}.$ Following the similar logic, the disorder parameter for the strong symmetry is $D_{1}$. Unlike for the weak symmetry, for the strong symmetry $O_{2}\propto\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|Z_{u,i}Z_{u,j}Z_{l,i}Z_{l,j}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩ is a valid order parameter. This is because $Z_{u,i}Z_{l,i}$ is charged under the strong symmetry (it fails to commute with $U_{u}$). $O_{1}$ also serves as an order parameter for the strong symmetry, so the strong symmetry is SSB if either $O_{2}$ or $O_{1}$ is nonzero (for a symmetry to be SSB, it suffices that any order parameter is nonzero).
@@
Strong and weak symmetries translate to different 1-form symmetries of the Choi state $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}$⟩, and the strong/weak order and disorder parameters map onto usual pure state order and disorder parameters of the Choi state.
@@
$U_{g,u}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\propto|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\qquad U_{g,l}^{T}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}\propto|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (22)
@@
$U_{g,u}U_{g,l}^{T}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}=|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},\quad g\in G.$ (23)
@@
$O_{2}=\lim\limits_{r\to\infty}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\rho|U_{h,u}U_{h,l}^{T}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}.$⟩ . (24)

### v3

@@ II.2.2 Relation to the Choi state
The last line can be expressed as a sum of expectation values in the unnormalized Choi state $|\tilde{\rho}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$.
@@
$|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=\frac{1}{\sqrt{\mathrm{Tr}(\rho^{2})}}\sum_{n}p_{n}|\psi_{n,u}\rangle\otimes|\psi_{n,l}^{*}\rangle,$ (14)
The unnormalized Choi state $|\tilde{\rho}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ is the same as $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ except with norm equal to the purity of $\rho$. If $\rho$ is a pure state, then $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ is a tensor product: $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=|\psi_{u}\rangle|\psi_{l}^{*}\rangle$. Therefore, a pure topologically ordered density matrix maps onto the topological order stacked with its time reversal partner. If $\rho$ is not pure, then there is generally some entanglement between spins in the upper and lower Hilbert spaces.
@@
$\mathrm{Tr}(\mathrm{Tr}_{C}(\rho)^{2})=\frac{1}{D_{C}}\sum_{\sigma_{n}\in C}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\tilde{\rho}|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\tilde{\rho}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$, (15)
@@
$\begin{split}&I_{2}(A:C|B)\\ &=-\log\frac{\sum_{\sigma_{n}\in C,\sigma_{m}\in A}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}}{\sum_{\sigma_{n}\in C,\sigma_{m}\in A}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}}\end{split}$ (16)
where we simplified the expression using the normalized Choi state $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=\frac{1}{\sqrt{\mathrm{Tr}(\rho^{2})}}|\tilde{\rho}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$. (LABEL:I2choi) is the main result of this section. It connects the Rényi-2 Markov length to connected correlation functions in the Choi state.
@@
$\begin{split}\Delta_{n,m}&=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\\ &-\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\end{split}$ (17)
@@
$\begin{split}&I_{2}(A:C|B)\\ &=-\log\left(1-\frac{\sum_{m,n}\Delta_{n,m}}{\sum_{\sigma_{n}\in C,\sigma_{m}\in A}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}}\right)\\ &\sim\frac{\sum_{m,n}\Delta_{n,m}}{\sum_{\sigma_{n}\in C,\sigma_{m}\in A}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}}\end{split}$ (18)
@@
where we expanded to leading order in $\frac{\sum_{m,n}\Delta_{n,m}}{\sum_{\sigma_{n}\in C,\sigma_{m}\in A}\mathopen{\hbox{\set@color${\langle}$}\kern-1.61806pt\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{n,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.61806pt\hbox{\set@color${\rangle}$}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.61806pt\hbox{\set@color${\langle}$}}\rho|\sigma_{m,u}\sigma_{m,l}^{\dagger}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.61806pt\hbox{\set@color${\rangle}$}}}$.
@@
Suppose that $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ displays exponential decay of connected correlations. Then $\Delta_{n,m}=0$ if $\sigma_{n,u}\sigma_{n,l}^{\dagger}=\mathbf{1}$ or $\sigma_{m,u}\sigma_{m,l}^{\dagger}=\mathbf{1}$ and $\Delta_{n,m}$ decays exponentially in $r$ otherwise. Therefore, $I_{2}(A:C|B)$ seems to also decays exponentially in $r$, and we obtain a finite Rényi-2 Markov length.
@@
For example, if $\rho=|\psi\rangle\langle\psi|$ is a pure state so $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=|\psi\rangle|\psi\rangle$, then assuming that $\rho$ can be written as a product of $|A\cup B\cup C|$ local projectors  [^13] Note that if $|\psi\rangle$ is related by a local unitary $U$ to a fixed point state $\prod_{i}P_{i}$ where each $P_{i}$ is a local projector, then $|\psi\rangle$ can also be written as a product of local projectors: $|\psi\rangle\langle\psi|=\prod_{i}U^{\dagger}P_{i}U$., the denominator of (LABEL:i2expand) is of order $2^{|A|+|C|}$. On the other hand, the numerator is bounded (very loosely) by $(2^{|A|+|C|}-2^{|A|}-2^{|C|})e^{-\mu r}$ for some constant $\mu$, so $I_{2}(A:C|B)$ decays exponentially with $r$. More generally when $\rho$ is mixed, then the denominator will typically be of order $e^{\nu(|A|+|C|)}$ for some $\nu$ (for example, if $\rho$ is the completely mixed state, $e^{\nu}=4$) and the numerator is loosely upper bounded by $(e^{\nu(|A|+|C|)}-e^{\nu|A|}-e^{\nu|C|})e^{-\mu r}$ so $I_{2}(A:C|B)$ is upper bounded by a function exponentially decaying in $r$. If instead $\Delta_{n,m}$ decays with a power law for $\mathcal{O}(e^{\nu(|A|+|C|)})$ operators, then $I_{2}(A:C|B)$ cannot be fit to an exponential decay. Attempting to fit $I_{2}(A:C|B)$ to exponential decay in $r$ would lead to a diverging length scale, which is the diverging Rényi-2 Markov length.
@@
Therefore, we expect that the Rényi-2 Markov length diverges whenever the Choi state encounters phase transitions where operators of the form $O_{u}O_{l}^{\dagger}$ have power-law decay of correlations. Note that since the expression (LABEL:I2choi) only has operators that are diagonal between the upper and lower Hilbert spaces, it is not bounded by decay of correlations of operators solely supported in a single Hilbert space, i.e. those of the form $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\sigma_{n,u}\sigma_{m,u}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$. However, due to the $\mathrm{SWAP}^{*}$ symmetry of the Choi state (explained further in Sec. II.4), the above scenario where there is only a diverging correlation length in $\mathcal{H}_{u}$ or $\mathcal{H}_{l}$ is ruled out.
@@
In the example of decohered toric code Fan et al. (2024); Bao et al. (2023); Lee et al. (2023); Zou et al. (2023); Sang and Hsieh (2024) (see Appendix D for a review) where we put the state $|tc\rangle\langle tc|$ through an incoherent $Z$ noise channel (4) and $X$ noise channel (same as (4) except with $Z_{e}$ replaced by $X_{e}$), the Choi state is initially described by two copies of toric code $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=|tc\rangle|tc\rangle$. Under decoherence, the Choi state generically demonstrates two transitions, because there is a transition from two copies of toric code to a single copy, and then from a single copy to the trivial state. In Appendix D.1, we show by explicit calculation in this example that $I_{2}(A:C|B)$ is related to the free energy of a point defect in the 2D classical Ising model, and $\xi_{2}$ diverges at two points.
@@ II.3 Strong and weak 1-form symmetries and order/disorder parameters
$U_{u}U_{l}^{T}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ (21)
where $u,l$ label the upper and lower Hilbert spaces corresponding to the ket and bra, respectively. The order parameter can be either two point correlator in the upper Hilbert space $\propto\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|Z_{u,i}Z_{u,j}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ or in the lower Hilbert space $\propto\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|Z_{l,i}Z_{l,j}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$, because neither $Z_{u,i}$ nor $Z_{u,j}$ commutes with $U_{u}U_{l}^{T}$, but both $Z_{u,i}Z_{u,j}$ and $Z_{l,i}Z_{l,j}$ commute with $U_{u}U_{l}^{T}$. Since only operators from one of the two Hilbert spaces is involved, we label the order parameter $O_{1}$ with subscript $1$. We cannot use $O_{2}$ as an order parameter because $Z_{u,i}Z_{l,i}$ is not charged under $U_{u}U_{l}^{T}$. By comparison, the disorder parameter is the symmetry action restricted to a subsystem (a $d$-ball in $d$ spatial dimensions). Since the weak symmetry acts on both Hilbert spaces, the disorder parameter lives on both Hilbert spaces as well, which is $\propto\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|\prod_{i\in A}X_{u,i}X_{l,i}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$, which explains the subscript $2$ in $D_{2}.$ Following the similar logic, the disorder parameter for the strong symmetry is $D_{1}$. Unlike for the weak symmetry, for the strong symmetry $O_{2}\propto\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|Z_{u,i}Z_{u,j}Z_{l,i}Z_{l,j}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ is a valid order parameter. This is because $Z_{u,i}Z_{l,i}$ is charged under the strong symmetry (it fails to commute with $U_{u}$). $O_{1}$ also serves as an order parameter for the strong symmetry, so the strong symmetry is SSB if either $O_{2}$ or $O_{1}$ is nonzero (for a symmetry to be SSB, it suffices that any order parameter is nonzero).
@@
Strong and weak symmetries translate to different 1-form symmetries of the Choi state $|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$, and the strong/weak order and disorder parameters map onto usual pure state order and disorder parameters of the Choi state.
@@
$U_{g,u}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\propto|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\qquad U_{g,l}^{T}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\propto|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}.$ (22)
@@
$U_{g,u}U_{g,l}^{T}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}},\quad g\in G.$ (23)
@@
$O_{2}=\lim\limits_{r\to\infty}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\rho|U_{h,u}U_{h,l}^{T}|\rho\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}.$ (24)